write a program on emu8086

Write a program that adds ten numbers and store the result in register AX,BX,CX,DX and DS



1
Expert's answer
2021-11-04T00:24:11-0400
start:

 

    mov ax,10

    add ax,2           ; ax = 10+2

    MOV DS,AX    ; move result to DS

               

    mov ax,3

    add ax,2           ; ax = 2+3

 

    mov bx,4

    add bx,8          ; bx = 4+8

               

    mov cx,4

    add cx,1    ; cx = 4+1

               

    mov dx,8

    add dx,4          ; dx = 8+4

               

    mov ah, 4ch

    int 21h

 

ends

end start ; set entry point and stop the assembler.